OpenStudy (anonymous):

(2x+3h)^2 i started off with 4x^2 now im stuck

OpenStudy (anonymous):

\[(2x+3h)^2=(2x+3h)(2x+3h)\] you have to do four multiplications

OpenStudy (anonymous):

first one does give \(4x^2\)

OpenStudy (anonymous):

16h^2

OpenStudy (anonymous):

\[(2x+3h)(2x+3h)=2x\times 2x+2x\times 3h+2x\times 3h+3h\times 3h\] is a start

OpenStudy (anonymous):

mostly you can skip some of that step, and go right to \[4x^2+6xh+6xh+9h^2\]

OpenStudy (anonymous):

4x^2+16h^2...i think im moving to fast and skipping steps

OpenStudy (anonymous):

four multiplications, just like if you were going to multiply \(23\times 23\)

OpenStudy (anonymous):

you have "middle terms" as well it is not just squaring the first and last term

OpenStudy (anonymous):

4x^2+12xh+9h^2 is what i got from your steps

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

how about \[23\times 23\] you would do 2 3 2 3 _____ 6 9 4 6 four multiplications, then add this is the same thing

OpenStudy (anonymous):

only this time you have variables so you get \(4x^2+6xh+6xh+9h^2\) or \(4x^2+12xh+9h^2\)

OpenStudy (anonymous):

ok so i see my mistake i forgot my middle terms and thats how you got 12xh